### v1

@@ Abstract
The support of artificial intelligence (AI) based decision-making is a key element in future 6G networks, where the concept of native AI will be introduced. Moreover, AI is widely employed in different critical applications such as autonomous driving and medical diagnosis. In such applications, using AI as black-box models is risky and challenging. Hence, it is crucial to understand and trust the decisions taken by these models. Tackling this issue can be achieved by developing explainable AI (XAI) schemes that aim to explain the logic behind the black-box model behavior, and thus, ensure its efficient and safe deployment. Recently, we proposed a novel perturbation-based XAI-CHEST framework that is oriented toward channel estimation in wireless communications. The core idea of the XAI-CHEST framework is to identify the relevant model inputs by inducing high noise on the irrelevant ones. This manuscript provides the detailed theoretical foundations of the XAI-CHEST framework. In particular, we derive the analytical expressions of the XAI-CHEST loss functions and the noise threshold fine-tuning optimization problem. Hence the designed XAI-CHEST delivers a smart input feature selection methodology that can further improve the overall performance while optimizing the architecture of the employed model. Simulation results show that the XAI-CHEST framework provides valid interpretations, where it offers an improved bit error rate performance while reducing the required computational complexity in comparison to the classical DL-based channel estimation.
@@
Generally speaking, the AI-enabled intelligent architecture for 6G networks defines several layers including the intelligent sensing layer [3]. It is worth mentioning that robust environment monitoring and data detection are of great interest in 6G smart applications like autonomous driving [4]. Note that ensuring the reliability of the intelligent sensing layer is highly impacted by the channel estimation accuracy since a precisely estimated channel response influences the follow-up equalization and decoding operations at the receiver, therefore, it affects the sensing accuracy [5]. In this context, channel estimation is one of the major physical (PHY) layer issues due to the doubly-selective nature of the channel in mobile applications. Conventional channel estimation schemes such as least squares (LS) and linear minimum mean square error (LMMSE) can not be practical in real-case scenarios. This is because the LS channel estimation scheme ignores the presence of noise in the estimation process and requires the transmission of a large number of pilots which decreases the transmission data rate. Whereas, the LMMSE channel estimator provides good performance assuming the prior knowledge of the channel and noise statistics in addition to its high computational complexity.
@@ I-D Motivation and Contributions
To the best of our knowledge, the methodology of deploying XAI schemes in PHY layer applications, specifically, channel estimation is still unclear. Noting that the proposed XAI-based schemes for network optimization [27], resource allocation [30], and secured IoT [33] can not be adapted to the PHY layer applications because in such applications there are no clear discriminative features within the model inputs. In this context, this paper aims to design a novel XAI framework for FNN-based channel estimation denoted as XAI-CHEST. This framework is based on a perturbation-based model-agnostic global pre-model methodology that jointly performs the channel estimation task and provides the corresponding interpretability. The XAI-CHEST concept has been partially proposed in [36], where the key idea is to provide the interpretability of black box models by inducing noise on the model input while preserving accuracy. The model inputs are then classified into relevant and irrelevant sets based on the induced noise. The contributions of this work can be summarized as follows:
@@ II-A OFDM transceiver
In this work, we use an orthogonal frequency division multiplexing (OFDM)-based transmission with non-linear radio frequency (RF) represented by the high power amplifier (HPA) at the OFDM transmitter. As shown in Figure 2, the first operation on the transmitter side is the binary bits generation. Generated bits are scrambled in order to randomize the bits pattern, which may contain long streams of s or s. The scrambled bits are then passed to the encoder, which introduces some redundancy into the bits stream. This redundancy is used for error correction that allows the receiver to combat the effects of the channel, hence reliable communications can be achieved.
@@ II-C DL-based Channel Estimation
We note that conventional STA channel estimation performs well in the low signal-to-noise ratio (SNR) region. However, it suffers from a considerable error floor in high SNR regions due to the large DPA demapping error resulting from (7) and the fixed frequency and time averaging coefficients in (9) and (10), respectively. Therefore, the conventional STA channel estimation scheme is not practical in real-case scenarios due to the high doubly-selective channel variations. As a workaround, a hidden layer FNN consisting of neurons per layer is utilized as a nonlinear post-processing unit following STA [8]. STA-FNN can better capture the time-frequency correlations of the channel samples, in addition to correcting the conventional STA estimation error.
@@ IV Simulation Results
This section illustrates the performance evaluation of the proposed XAI-CHEST framework, where BER performance of STA-FNN channel estimation scheme is analyzed taking into consideration full, relevant, and irrelevant subcarriers. First of all, we start with the noise weight threshold fine-tuning, where the simulation-based solution of (17) is provided. After that, the performance evaluation is performed according to several criteria including the (i) modulation order, (ii) frequency selectivity of the channel, (iii) training SNR, and (iv) conventional channel estimation accuracy. Finally, a detailed computational complexity analysis is discussed where we show that further significant reduction in the overall computational complexity can be achieved by employing only the relevant subcarriers identified by the proposed XAI-CHEST framework.
We note that the considered channel models [41] as shown in Table I: (i) Low-frequency selectivity (LFS), where VTV Expressway (VTV-EX) scenario is employed. (ii) High-frequency selectivity (HFS), where VTV Expressway Same Direction with Wall (VTV-SDWW) scenario is considered. In both scenarios, Doppler frequency Hz is considered. Both the and models are trained using a OFDM symbols dataset, splitted into training, and testing. ADAM optimizer is used with a learning rate with batch size equals for epoch. Simulation parameters are based on the IEEE 802.11p standard [5], where the comb pilot allocation is used so that , , , and . Table II summarizes the simulation parameters considered in this work. Finally, we note that the STA channel estimation is considered as an initial estimation prior to the FNN processing. Hence, , unless stated otherwise.
@@ IV-B Impact of modulation order
However, we can notice that for higher the modulation order, the number of neglected subcarriers increases. This is because the conventional STA channel estimates at these subcarriers are so noisy, so the model neglects them. It is worth mentioning that, the model neglects subcarriers due to two main reasons: (i) LFS: The channel variation among the subcarriers is slow, so few subcarriers are required to accurately estimate the channel, as we will discuss in the next Section. (i) HFS: Here the channel variation is significant among the subcarriers, thus, the model should consider more relevant subcarriers to guarantee good channel estimation accuracy. However, this is subject to the condition where the conventional estimated channel at the considered subcarriers is useful and not so noisy. Therefore, in the HFS channel model, more relevant subcarriers are needed and this is shown in generally shifting the noise weight distribution towards zero. However, for higher modulation orders, mainly 64QAM, the neglected subcarriers are huge due to the bad channel estimation quality at these subcarriers. Hence, avoiding them is useful to guarantee BER performance. We note that the four pilot subcarriers are assigned the lowest noise weight for all the modulation orders. Therefore, the model is able to classify the pilots as the most relevant subcarriers regardless of the channel’s high selectivity and the employed modulation order.
@@ IV-E Impact of training SNR
The sensitivity of the model training considering different SNR values is analyzed in this section. Figure 11(a) shows the noise distribution when considering several training SNRs employing the LFS channel model and QPSK modulation order. Starting by training SNR = dB, we can see that the pilot subcarriers are assigned noise weight and the distribution is flattened along the entire range. This reveals that even though the pilots have accurate channel estimates, due to the dominant impact of AWGN noise, the model is not able to assign the lowest noise weight to the pilot subcarriers. It is worth mentioning that when training on SNR = dB, the model starts to identify the pilot subcarriers as the most relevant subcarriers by assigning to two pilots the lowest noise weight, i.e., . Moreover, as the training SNR increases, the noise distribution is shifted more towards one, signifying that the model is better identifying the relevant and irrelevant subcarriers. Figure 11(b) shows the BER performance when the model is trained on a specific SNR and tested on the entire SNR range. We can notice that training on higher SNR gives better performance than training on the lower SNR due to the fact the AWGN noise is negligible at high SNRs, thus the model can learn more efficiently the channel. In addition, the trained model on high SNR can perform well when tested on lower SNRs due to the generalization ability of FNN networks. In conclusion, training on low SNR values leads to a limited performance improvement over the conventional STA channel estimation. Whereas training on high SNR allows the smart feature selection resulting in optimizing the model input size, as well as significantly improving the BER performance in comparison to the conventional STA channel estimation.
@@ IV-F Impact of conventional channel estimation accuracy
To further analyze the impact of the conventional channel estimation which is implemented prior to the FNN processing on the noise weight distribution, we considered in this section the DPA-FNN [7] and TRFI-FNN [9] channel estimation schemes in addition to STA-FNN. We note that we consider the HFS channel model with QPSK modulation in this analysis since it is more challenging.
@@ IV-G Computational complexity reduction
V Conclusion
Ensuring the transparency and trustworthiness of AI is a crucial need for its efficient and safe deployment in critical applications. In this paper we designed a novel XAI-CHEST framework that provides the interpretability of the FNN models employed in the channel estimation application. The XAI-CHEST framework aims to classify the black-box model inputs into relevant and irrelevant inputs by using a perturbation-based methodology. We developed the theoretical foundations of the XAI-CHEST framework by formalizing the related loss functions. Moreover, the noise threshold fine-tuning optimization problem has been analytically derived. Extensive simulations have been conducted where the results reveal that a trustworthy, optimized, and low-complexity channel estimation scheme can be designed by selecting only the relevant inputs. As a future perspective, the gradient-based XAI for wireless communication will be investigated, where we will study the possibility of providing robust interpretation by using the internal architecture of the FNN model.

### v2

@@ Abstract
The support of artificial intelligence (AI) based decision-making is a key element in future 6G networks. Moreover, AI is widely employed in critical applications such as autonomous driving and medical diagnosis. In such applications, using AI as black-box models is risky and challenging. Hence, it is crucial to understand and trust the decisions taken by these models. Tackling this issue can be achieved by developing explainable AI (XAI) schemes that aim to explain the logic behind the black-box model behavior, and thus, ensure its efficient and safe deployment. Recently, we proposed a novel perturbation-based XAI-CHEST framework that is oriented toward channel estimation in wireless communications. The core idea of the XAI-CHEST framework is to identify the relevant model inputs by inducing high noise on the irrelevant ones. This manuscript provides the detailed theoretical foundations of the XAI-CHEST framework. In particular, we derive the analytical expressions of the XAI-CHEST loss functions and the noise threshold fine-tuning optimization problem. Hence the designed XAI-CHEST delivers a smart input feature selection methodology that can further improve the overall performance while optimizing the architecture of the employed model. Simulation results show that the XAI-CHEST framework provides valid interpretations, where it offers an improved bit error rate performance while reducing the required computational complexity in comparison to the classical DL-based channel estimation.
@@
Generally speaking, the AI-enabled intelligent architecture for 6G networks defines several layers including the intelligent sensing layer [3]. It is worth mentioning that robust environment monitoring and data detection are of great interest in 6G smart applications like autonomous driving [4]. Note that ensuring the reliability of the intelligent sensing layer is highly impacted by the channel estimation accuracy since a precisely estimated channel response influences the follow-up equalization and decoding operations at the receiver, therefore, it affects the sensing accuracy [5]. In this context, channel estimation is one of the major physical (PHY) layer issues due to the doubly-selective nature of the channel in mobile applications. Conventional channel estimation schemes such as least squares (LS) ignores the presence of noise in the estimation process and requires the transmission of a large number of pilots which decreases the transmission data rate. Whereas, the linear minimum mean square error (LMMSE) channel estimator provides good performance assuming the prior knowledge of the channel and noise statistics in addition to its high computational complexity.
@@ I-D MOTIVATION AND CONTRIBUTIONS
To the best of our knowledge, the methodology of deploying XAI schemes in PHY layer applications, specifically, channel estimation is still unclear. Noting that the proposed XAI-based schemes for network optimization [27], resource allocation [30], and secured IoT [33] can not be adapted to the PHY layer applications because in such applications there are no clear discriminative features within the model inputs. In this context, this paper aims to design a novel XAI framework for FNN-based channel estimation, denoted as XAI-CHEST. This framework is based on a perturbation-based model-agnostic global pre-model methodology that jointly performs the channel estimation task and provides the corresponding interpretability. The XAI-CHEST concept has been partially proposed in [36], where the key idea is to provide the interpretability of black box models by inducing noise on the model input while preserving accuracy. The model inputs are then classified into relevant and irrelevant sets based on the induced noise. The contributions of this work can be summarized as follows:
@@ II-A OFDM TRANSCEIVER
In this work, we consider single-input and single-output (SISO) orthogonal frequency division multiplexing (OFDM)-based transmission with non-linear radio frequency (RF) represented by the high power amplifier (HPA) at the OFDM transmitter. As shown in Figure 2, the first operation on the transmitter side is the binary bits generation. Generated bits are scrambled in order to randomize the bits pattern, which may contain long streams of s or s. The scrambled bits are then passed to the encoder, which introduces some redundancy into the bits stream. This redundancy is used for error correction that allows the receiver to combat the effects of the channel, hence reliable communications can be achieved.
@@ II-C DL-BASED CHANNEL ESTIMATION
We note that conventional STA channel estimation performs well in the low signal-to-noise ratio (SNR) region. However, it suffers from a considerable error floor in high SNR regions due to the large DPA demapping error resulting from (7) and the fixed frequency and time averaging coefficients in (9) and (10), respectively. Therefore, the conventional STA channel estimation scheme is not practical in real-case scenarios due to the high doubly-selective channel variations. As a workaround, a hidden layer FNN consisting of neurons per layer is utilized as a nonlinear post-processing unit following STA. As shown in [8], STA-FNN can better capture the frequency correlations of the channel samples, in addition to correcting the conventional STA estimation error.
@@ IV SIMULATION RESULTS
This section illustrates the performance evaluation of the proposed XAI-CHEST framework, where BER performance of STA-FNN channel estimation scheme is analyzed taking into consideration full, relevant, and irrelevant subcarriers. First of all, we start with the noise weight threshold fine-tuning, where the simulation-based solution of (17) is provided. After that, the performance evaluation is performed according to several criteria including the (i) modulation order, (ii) frequency selectivity of the channel, (iii) training SNR, and (iv) conventional channel estimation accuracy. Finally, a detailed computational complexity analysis is discussed where we show that further significant reduction in the overall computational complexity can be achieved by employing only the relevant subcarriers identified by the proposed XAI-CHEST framework. We note that the considered channel models [41] are shown in Table I: (i) Low-frequency selectivity (LFS), where VTV Expressway (VTV-EX) scenario is employed. (ii) High-frequency selectivity (HFS), where VTV Expressway Same Direction with Wall (VTV-SDWW) scenario is considered. In both scenarios, Doppler frequency Hz is considered. Both the and models are trained using a OFDM symbols dataset, splitted into training, and testing. ADAM optimizer is used with a learning rate with batch size equals for epoch. Simulation parameters are based on the IEEE 802.11p standard [5], where the comb pilot allocation is used so that , , , and . Table II summarizes the simulation parameters considered in this work. Finally, we note that the STA channel estimation is considered as an initial estimation prior to the FNN processing. Hence, , unless stated otherwise.
@@ IV-B IMPACT OF MODULATION ORDER
However, we can notice that for higher the modulation order, the number of neglected subcarriers increases. This is because the conventional STA channel estimates at these subcarriers are so noisy, so the STA-FNN model neglects them. It is worth mentioning that, the STA-FNN model treats the conventional STA estimated channels separately and does not consider the time correlation between successive OFDM symbols due to the architectural design of the FNN network. Hence, the channel tracking over time is applied within the conventional STA scheme, where the FNN model is used to capture the frequency correlation of the channel samples as well as coping with the conventional STA estimation error. In this context, the STA-FNN model neglects subcarriers due to two main reasons: (i) LFS: The channel variation among the subcarriers is slow, so few subcarriers are required to accurately estimate the channel, as we will discuss in the next Section. (i) HFS: Here the channel variation is significant among the subcarriers, thus, the model should consider more relevant subcarriers to guarantee good channel estimation accuracy. However, this is subject to the condition where the conventional estimated channel at the considered subcarriers is useful and not so noisy. Therefore, in the HFS channel model, more relevant subcarriers are needed and this is shown in generally shifting the noise weight distribution towards zero. However, for higher modulation orders, mainly 64QAM, the neglected subcarriers are huge due to the bad channel estimation quality at these subcarriers. Hence, avoiding them is useful to guarantee BER performance. We note that the four pilot subcarriers are assigned the lowest noise weight for all the modulation orders. Therefore, the model is able to classify the pilots as the most relevant subcarriers regardless of the channel’s high selectivity and the employed modulation order.
@@ IV-E IMPACT OF TRAINING SNR
The sensitivity of the model training, considering different SNR values, is analyzed in this section. Figure 11(a) shows the noise distribution when considering several training SNRs employing the LFS channel model and QPSK modulation order. Starting by training SNR = dB, we can see that the pilot subcarriers are assigned noise weight and the distribution is flattened along the entire range. This reveals that even though the pilots have accurate channel estimates, due to the dominant impact of AWGN noise, the model is not able to assign the lowest noise weight to the pilot subcarriers. It is worth mentioning that when training on SNR = dB, the model starts to identify the pilot subcarriers as the most relevant subcarriers by assigning to two pilots the lowest noise weight, i.e., . Moreover, as the training SNR increases, the noise distribution is shifted more towards one, signifying that the model is better identifying the relevant and irrelevant subcarriers. Figure 11(b) shows the BER performance when the model is trained on a specific SNR and tested on the entire SNR range. We can notice that training on higher SNR gives better performance than training on the lower SNR due to the fact the AWGN noise is negligible at high SNRs, thus the model can learn more efficiently the channel. In addition, the trained model on high SNR can perform well when tested on lower SNRs due to the generalization ability of FNN networks. In conclusion, training on low SNR values leads to a limited performance improvement over the conventional STA channel estimation. Whereas training on high SNR allows the smart feature selection resulting in optimizing the model input size, as well as significantly improving the BER performance in comparison to the conventional STA channel estimation.
@@ IV-F IMPACT OF CONVENTIONAL CHANNEL ESTIMATION ACCURACY
To further analyze the impact of the conventional channel estimation, which is implemented prior to the FNN processing on the noise weight distribution, we considered in this section the DPA-FNN [7] and TRFI-FNN [9] channel estimation schemes in addition to STA-FNN. We note that we consider the HFS channel model with QPSK modulation in this analysis since it is more challenging.
@@ IV-G COMPUTATIONAL COMPLEXITY REDUCTION
V CONCLUSION AND FUTURE PERSPECTIVES
Ensuring the transparency and trustworthiness of AI is a crucial need for its efficient and safe deployment in critical applications. In this paper we designed a novel XAI-CHEST framework that provides the interpretability of the FNN models employed in the channel estimation application. The XAI-CHEST framework aims to classify the black-box model inputs into relevant and irrelevant inputs by using a perturbation-based methodology. We developed the theoretical foundations of the XAI-CHEST framework by formalizing the related loss functions. Moreover, the noise threshold fine-tuning optimization problem has been analytically derived. Extensive simulations have been conducted where the results reveal that a trustworthy, optimized, and low-complexity channel estimation scheme can be designed by selecting only the relevant inputs.
As a future perspective, three main research directions could be established:
XAI-CHEST for RNN-based channel estimation: The functionality of the proposed XAI-CHEST framework is limited since it is not yet adapted to cope with the time variation of the wireless channel. Hence, it is essential to extend the XAI-CHEST framework to deal with RNN-based channel estimation. Thanks to the RNN memory that allows the prediction of the current channel based on previously estimated channels, RNN networks such as long short-term memory (LSTM) and gated recurrent unit (GRU) are able to perform channel estimation and tracking over time. Therefore, adapting the XAI-CHEST to the RNN-based channel estimation provides better noise allocation that varies over time among the received OFDM symbols. Moreover, investigating the impact of the RNN memory size of the efficiency of the noise allocation could provide further performance-complexity trade-offs.
XAI-CHEST for MIMO-OFDM: The combination of multi-antenna and multi-carrier technologies is a promising technique for ensuring the efficiency of high-speed transmission in wireless communication systems. However, providing performance-complexity trade-offs is crucial for better designing the MIMO-OFDM receiver. In this context, extending the proposed XAI-CHEST for MIMO-OFDM could be beneficial in adapting the size of the considered MIMO system based on the channel correlation and the desired performance.
Gradient-assisted XAI-CHEST framework: The proposed XAI-CHEST framework is based on an external model-agnostic perturbation-based methodology. Hence, the provided model’s interpretability is impacted only by studying the influence of the model inputs on its decision, where the internal architecture of the model remains black-box. Therefore, the current XAI-CHEST framework provides a smart input filtering strategy where model-driven optimization is still unexplored. In this context, investigating internal gradient-based XAI schemes [44] and integrating them within the XAI-CHEST framework will provide a double optimization strategy that leads to filtering the relevant model inputs, as well as fine-tuning the model architecture where relevant layers and neurons are preserved.